The sum of three angles is 120 degrees. The second angle is 5 degrees more than the first. The third angle is three times the second angle. Find the measure of each angle.
Answer by rfer(16322)

x+x+5+3(x+5)=120
x+x+5+3x+15=120
5x=120-20
x=100/5
x=20
x+5=25
3(x+5)=75